I am trying to solve this quadratic growth question. A town experiences quadratic growth of the form (P=7t^2 + Po) where (t) is the time in years from 2010. (Po) is the population , and I have determined that to be 7806. How do I make this into a quadratic equation and then factor it? thank you in advance for your help?
Answer by stanbon(75887)

A town experiences quadratic growth of the form (P=7t^2 + Po) where (t) is the time in years from 2010.
(Po) is the population , and I have determined that to be 7806.
How do I make this into a quadratic equation and then factor it?
thank you in advance for your help?
P(t) = 7t^2+7806
Factoring:
7t^2+7806 =0
t^2 = (7806/7)
t = sqrt(7806/7)
t = +/-33.3937
P(t) = 7(t-33.3937)(t+33.3937)
=====================
Cheers,
Stan H.